OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (anonymous):

use y-y1 = m(x-x1) where m = slope and x1,y1 is the coordinates of the point standard form is ax + by = c